### v1

# Profiling stellar environments of gravitational wave sources

Abstract
Gravitational waves (GWs) have enabled direct detections of compact binary coalescences (CBCs). However, their poor sky localisation and the typical lack of observable electromagnetic (EM) counterparts make it difficult to confidently identify their hosts, and study the environments that nurture their evolution. In this work, we show that detailed information of the host environment (e.g. the mass and steepness of the host potential) can be directly inferred by measuring the kinematic parameters (acceleration and its time-derivatives) of the binary’s center of mass using GWs alone, without requiring an EM counterpart. We consider CBCs in various realistic environments such as globular clusters, nuclear star clusters, and active galactic nuclei disks to demonstrate how orbit and environment parameters can be extracted for CBCs detectable by ground- and space-based observatories, including the LIGO detector at A+ sensitivity, Einstein Telescope of the XG network, LISA, and DECIGO, on a single-event basis. These constraints on host stellar environments promise to shed light on our understanding of how CBCs form, evolve, and merge.
Introduction.— A number of efforts trying to probe the provenance of compact binary coalescences (CBCs) detected with the LIGO-Virgo-KAGRA (LVK) gravitational wave (GW) detector network Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021); Abbott et al. (2023) are currently underway (see, e.g., Mapelli (2021) for a review). Indeed, determining the environment in which the CBC formed, evolved, and merged is highly non-trivial for two primary reasons. First, the GW waveform pertaining to the CBC does not always contain signatures of the CBC’s formation channel. While statistical arguments based on estimates of the CBC’s intrinsic source parameters could provide hints of its host environment, these cannot in general conclusively rule out all other possible environments in favor of the statistically preferred one Zevin et al. (2021); Pierra et al. (2024). Second, the lack of observed electromagnetic counterparts for all but one of the detected CBC events Abbott et al. (2023), in tandem with poor localisation sky-areas spanning $\mathcal{O}(10)$ sq. deg. or larger, make it difficult to identify the host via direct observations Chen and Holz (2016). The inclusion of additional detectors in current and forthcoming observing runs, such as KAGRA and LIGO-India Saleem et al. (2022), as well as planned and proposed next-generation ground and space-based missions, will have markedly improved localisation sky-areas, even going below $\mathcal{O}(1)$ sq. deg. for space-based missions Zhang et al. (2021). However, confident identification of the CBC’s host will require arc- or sub-arc-second localisations, which will in general not be achievable even by future detectors, especially at large redshifts.
Kinematic parameters (such as acceleration) associated with the motion of the center of mass (CoM) of the CBC, could serve as messengers of the CBC’s environment. Past work has shown that a finite line-of-sight acceleration of the CBC’s CoM will modulate the GW waveform with respect to one produced by a non-accelerated CBC  Vijaykumar et al. (2023); Bonvin et al. (2017); Yunes et al. (2011). At leading order, these modulations occur at $-4$ PN (Post Newtonian order). Higher time derivatives of this acceleration (e.g. jerk, snap, etc) cause modulations that start at even lower PN orders (e.g. $-8$ PN, $-16$ PN, etc) Yunes et al. (2011). Thus, detectors with sensitivities at low frequencies, such as space-based ones, are best suited to extract these kinematic parameters, although sufficiently large accelerations might be observable even with ground-based detectors Vijaykumar et al. (2023).
Most proposed formation channels of stellar-mass CBCs suggest that these either underwent isolated evolution in the galactic field, or were assembled dynamically in dense stellar environments such as globular clusters (GCs), nuclear star clusters (NSCs), and active galactic nuclei (AGNs). These environments have characteristic properties, in particular, mass and potential profiles, which govern the kinematics of the CoMs of the CBCs they host, thereby imprinting themselves on the emanated GWs. In this Letter, we demonstrate how these properties can be directly extracted from the CBC’s GWs, thus effectively identifying the CBC’s environment on a single-event basis.
We calculate the corrections to the GW phase as a function of the kinematic parameters. We then evaluate the relationship between the kinematic parameters and the properties of the environment (e.g. steepness of the potential), as well as the orbit of the CoM of the binary. We are thus able to extract and constrain the corresponding parameters and explore the precision of their inference in various single-detector configurations, viz. a LIGO detector at A+ sensitivity Abbott et al. (2020), Einstein Telescope (ET) Punturo et al. (2010) of the XG network Reitze et al. (2019); Punturo et al. (2010), the LISA Amaro-Seoane et al. (2017) and DECIGO Sato et al. (2017) space-based detectors.
Method.— The total mass of a binary $M=m_{1}+m_{2}$ in the detector frame as inferred from the GW signal is related to the source frame total mass $M_{\mathrm{src}}$ as:
[…]
where $z_{\mathrm{cos}}$ is the cosmological redshift, and $z_{\mathrm{dop}}$ is the Doppler shift caused by any line of sight velocity (LOSV) of the CoM of the binary [^1] From hereon, $M$ will denote the detector frame total mass.. A constant LOSV is fully degenerate with the source frame mass. However, a time-varying LOSV results in a time-varying Doppler shift and induces a time-varying detector frame mass, allowing time derivatives of the LOSV to be measured. Modulations to the GW waveform appear at different Post-Newtonian (PN) orders depending on the order of the time derivative of LOSV causing the modulation.
For a CoM with a finite LOS acceleration (LOSA), and higher LOSV ($v_{l}$) time derivatives, the detector frame mass becomes:
$M_{z}=M\left(1+\sum_{n=1}^{\infty}\Gamma_{n}(t_{o}-t_{c})^{n}\right)$ (2)
where $\Gamma_{n}\equiv\left.\frac{1}{c}\frac{1}{n!}\frac{d^{n}v_{l}}{dt_{o}^{n}}\right|_{t_{o}=t_{c}}$, $t_{o}$ is the observation time, $t_{c}$ is the time at coalescence, and $c$ is the speed of light. Note that Eq. (2) is valid under the assumption that $|\Gamma_{n}(t_{o}-t_{c})^{n}|$ are $\ll$ 1222As a consistency check of this condition, we ensure that $|\Gamma_{n}(t_{o}-t_{c})^{n}|\leq 0.05$., where special relativistic effects may be neglected.
The corresponding modulated GW waveform, in the stationary phase approximation, may be written as Buonanno et al. (2009):
$\tilde{h}(f)=\mathcal{A}f^{-7/6}e^{i\left(\Psi(f)+\Delta\Psi_{-4n}(f)\right)}$ (3)
where $\mathcal{A}\propto\frac{\mathcal{M}^{5/6}}{D_{L}}$ is the amplitude, $\mathcal{M}$ is the chirp mass, $D_{L}$ is the luminosity distance, $\Psi(f)$ is the unperturbed phase given by Eq. 3.18 of Buonanno et al. (2009), and $\Delta\Psi_{-4n}(f)$ is a $-4n$ PN phase correction due to $n^{th}$ derivative of $v_{l}$. At leading order, this is given by Tiwari et al. (2024a):
$\Delta\Psi_{-4n}(f)=-\frac{2^{-8n-7}5^{n+1}v^{-8n-5}\left(-\frac{M}{\eta}\right)^{n}}{\eta(n+1)}\Gamma_{n}$ (4)
where $v=(\pi Mf)^{1/3}$, $\eta\equiv m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is the symmetric mass ratio, and $f$ is the observed GW frequency.
Positive powers of a kinematic parameter can contribute to PN orders lower than the corresponding leading order. For example, the $-8$PN phase correction $\Delta\Psi_{-8}(f)$ will have contributions from the line of sight jerk (LOSJ) $\Gamma_{2}$ of the CoM, as well as the LOSA term proportional to $\Gamma_{1}^{2}$. We account for these corrections as well, down to $-24$ PN, below which we do not expect PN terms to be well constrained even with detectors sensitive down to mili-hertz frequencies [^3] See Figure 6 and the corresponding section in the Supplemental Material Tiwari et al. (2024a)..
Figure 1: A schematic representation of a binary orbiting around an SMBH (or, any center of potential) in a circular orbit. Here, $\iota_{\mathrm{out}}$ is the angle between our LOS $\bm{n}$ and the angular momentum vector of the outer orbit, which is along $\bm{e}_{z}$.
Figure 2: Circular outer-orbit around an SMBH: The precision of the extraction of parameters over a grid of SMBH masses $M_{\rm SMBH}$ and radii of outer orbit $R$ for a binary system of component masses (from left to right): $5-1.4\,M_{\odot}$ at $\mathrm{100Mpc}$ in A+, $30-30\,M_{\odot}$ at $100\mathrm{Mpc}$ in ET, and $100-100\,M_{\odot}$ at $1\mathrm{Gpc}$ in DECIGO and LISA, with $\cos\vartheta=0.9$. Top Panels: Relative error in the extraction of mass of SMBHs. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm SMBH}}{M_{\rm SMBH}}$. Bottom Panel: Relative error in the extraction of radii of the outer orbit. The ‘dashed’ lines are the lines of constant $\frac{\Delta R}{R}$. The pink regions correspond to the parameter space where either the relative errors are greater than 1 or we do not sample them due to the non-measurability of the kinematic parameters. The blue lines demarcate the regions (lower left) where we do not perform the Fisher matrix analysis, for, in these regions, the assumption $|\Gamma_{n}(t_{o}-t_{c})^{n}|\ll 1$ is no longer valid. Note that the arrows point towards the region of increasing precision in the inference of parameters.
Figure 3: Eccentric outer-orbit around an SMBH: The precision of the extraction of parameters over a grid of SMBH masses $M_{\rm SMBH}$ and semi-major axes $R$ of the outer orbit for a binary system of component masses (from left to right) considered in Figure 2 with $e=0.5$ and $x\approx 1.25$. $x$ was drawn randomly from a uniform distribution between $\frac{1}{1-e}$ and $\frac{1}{1+e}$. Top Panel: Relative error in the extraction of mass of SMBHs. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm SMBH}}{M_{\rm SMBH}}$. Bottom Panel: Relative error in the extraction of the location of binary in the outer orbit. The ‘dotted’ lines are the lines of constant $\frac{\Delta x}{x}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
We consider three environments that could impart time-varying LOSVs to a CBC: A) the vicinity of a supermassive black hole (SMBH) as a proxy for an AGN environment, B) a globular cluster (GC), and C) a nuclear star cluster (NSC). To analyse the kinematics of the CoM of the CBC in these environments, we set up a Cartesian coordinate system in Figure 1 placed at the center of the gravitating source. Let $\iota_{\mathrm{out}}$ be the angle between our LOS and the angular momentum vector of the outer orbit, and let the projection vector of our LOS onto the outer orbital plane be along $\bm{e}_{y}$. Then:
$\bm{n}$ $=\sin\iota_{\mathrm{out}}\bm{e}_{y}+\cos\iota_{\mathrm{out}}\bm{e}_{z}$ (5)
$\bm{e}_{r}(\vartheta)$ $=\cos\vartheta\bm{e}_{x}+\sin\vartheta\bm{e}_{y}$
$\bm{e}_{\vartheta}(\vartheta)$ $=-\sin\vartheta\bm{e}_{x}+\cos\vartheta\bm{e}_{y}$
where $\vartheta$ is the angular location of the binary in the orbital plane with respect to the x-axis, and $r$ is the radial location.
Let us further define a vector $\bm{\mathcal{V}}(r,\vartheta)$ such that:
$\bm{\mathcal{V}}(r,\vartheta)\equiv\mathcal{V}_{r}(r)\bm{e}_{r}(\vartheta)+\mathcal{V}_{\vartheta}(r)\bm{e}_{\vartheta}(\vartheta)$ (6)
where $r\equiv r(t)$, $\vartheta\equiv\vartheta(t)$, and $\bm{e}_{r}(\vartheta)$ and $\bm{e}_{\vartheta}(\vartheta)$ are the unit vectors. Then observing that $\dot{\bm{e}_{r}}=\dot{\vartheta}\bm{e}_{\vartheta}$ and $\dot{\bm{e}_{\vartheta}}=-\dot{\vartheta}\bm{e}_{r}$, we can write:
$\dot{\bm{\mathcal{V}}}(r,\vartheta)=(\mathcal{V}^{\prime}_{r}\dot{r}-\mathcal{V}_{\vartheta}\dot{\vartheta})\bm{e}_{r}+(\mathcal{V}^{\prime}_{\vartheta}\dot{r}+\mathcal{V}_{r}\dot{\vartheta})\bm{e}_{\vartheta}$ (7)
where $\cdot\equiv\frac{d}{dt}$ and ${}^{\prime}\equiv\frac{d}{dr}$. Using this relation iteratively, we calculate different kinematic parameters of the CoM of the binary and then project them onto our LOS ($\bm{n}$). This LOS component of the kinematic parameter will be a function of environment and orbit parameters, such as the steepness of the gravitational potential profile, and the location of the CBC with respect to the center of the potential, at coalescence [^4] For instance, a circular orbit of radius $R$ in a Keplerian potential around a SMBH of mass $M_{\rm SMBH}$ would have $\Gamma_{n}\sim\Omega^{n+1}R$, where $\Omega^{2}\sim M_{\rm SMBH}/R^{3}$.. We provide the derivation as well as the mathematical expressions of the LOS components in the Supplementary Material Tiwari et al. (2024a).
To assess our ability to constrain these parameters in various observing scenarios, we adopt a Fisher Matrix approach Cutler and Flanagan (1994). The PN corrections to the GW phase due to the kinematic parameters are rewritten in terms of the environment and outer-orbit parameters. A Fisher Matrix on these is constructed and then inverted to yield a Covariance Matrix, which provides the Gaussian approximation to the GW likelihood. The environment parameters are sampled from this likelihood using emcee Foreman-Mackey et al. (2013), and the relative r.m.s errors on these parameters are estimated. We delineate these steps in detail in the Supplementary Material Tiwari et al. (2024a).
Results.— We present results for three CBC environments: vicinity of SMBH (Keplerian potential), GC (Plummer Plummer (1911) profile), NSC (Bahcall-Wolf Hoang et al. (2018) profile); three orbital shapes: circular, elliptic and effective Keplerian [^5] What we can call an “effective Keplerian orbit” is a bound orbit that is not necessarily closed.; and four single-detector configurations: LIGO at A+ sensitivity, Einstein Telescope (ET), DECIGO, and LISA. While circular orbits are possible in all attractive potentials, closed elliptical orbits are only possible in attractive potentials with $1/r$ and $1/r^{2}$ profiles. Moreover, only those orbits with a changing $r$ will be sensitive to the steepness of the potential profile. Thus, for Bahcall-Wolf and Plummer potentials, we only consider effective Keplerian orbits by evolving both $r$ and $\vartheta$ over the duration of the GW signal within the frequency band of the detectors.
For binaries in a circular orbit around an SMBH with Schwarzschild radius $R_{\rm s}$, the kinematic parameters along our LOS depend on the mass of the SMBH ($M_{\rm SMBH}$), the radius of the circular orbit ($R$), and $\cos\vartheta$ (see Figure 1). We use the first three derivatives of LOSV, viz. LOSA, LOSJ, and LOS snap (LOSS), to uniquely determine these quantities [^6] Note that $M_{\rm SMBH}$ and $R$ will still be degenerate with $\iota_{\mathrm{out}}$ and this degeneracy can not be broken even if we consider one more derivative. Similar degeneracies exist between the environment/orbit parameters, and $\iota_{\mathrm{out}}$, for all scenarios considered in this work. We set $\sin\iota_{\mathrm{out}}=1$ throughout; thus, the relative errors in the inference of environment and orbit parameters should be thought of as lower limits..
Figure 2 shows the relative errors in the measurements of $M_{\rm SMBH}$ (top panel) and $R$ (bottom panel) over a $M_{\rm SMBH}-R$ grid. We denote the inference of a parameter as “successful” if the relative r.m.s. error is less than unity. We find that:
  1. 1.
    At A+ sensitivity, we can successfully extract the mass of a $M_{\mathrm{SMBH}}=10^{5}\,M_{\odot}$ SMBH and localise a $5-1.4\,M_{\odot}$ binary having optimal SNR [^7] Note that for all of the cases considered in this work, the SNRs quoted are for face-on inner orbits. $\sim 103$ corresponding to a luminosity distance $D_{L}=100\,{\rm Mpc}$, out to $R\sim 200\,R_{s}$.
  2. 2.
    With ET, we can successfully extract the mass of the same SMBH and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 6610$ corresponding to $D_{L}=100\,{\rm Mpc}$, out to $R\sim 250\,R_{s}$, and within a precision of $\sim 10\%$ up to $R\sim 100\,R_{s}$.
  3. 3.
    We can successfully infer the mass of an SMBH as large as $10^{8}\,M_{\odot}$, and localise a $100-100\,M_{\odot}$ binary having optimal SNR $\sim 7624\;(10)$ corresponding to $D_{L}=1{\rm Gpc}$, out to $\sim 10^{5}\,R_{s}$ ($\sim 2\times 10^{4}\,R_{s}$) with DECIGO (LISA).
Furthermore, for tight orbits around low mass black holes we can extract these quantities with sub-percent precision. All results in this analysis assume $\cos\vartheta$ to be 0.9. In the Supplemental Material Tiwari et al. (2024a), we show that the precision of the inferred environment and orbit parameters do not depend strongly on the value of $\cos\vartheta$.
For binaries in an elliptical orbit around an SMBH, the kinematic parameters along our LOS depend on $M_{\rm SMBH}$, semi-major axis ($R$), eccentricity ($e$), and $x\equiv R/r$ where $r$ is distance of the CoM of the binary from the gravitating center, at $t_{c}$. Therefore, together with LOSA, LOSJ, and LOSS, we will need one more derivative of LOSV: LOS crackle (LOSC), to uniquely determine these quantities (see Tiwari et al. (2024a) for the expressions of kinematic parameters).
Figure 3 shows the relative errors in the measurements of $M_{\rm SMBH}$ (top panel) and $x$ (bottom panel) over a $M_{\rm SMBH}-R$ grid. We find the following:
  1. 1.
    With LIGO at A+ sensitivity, we can successfully extract the mass of a $10^{5}\,M_{\odot}$ SMBH and localise a $5-1.4\,M_{\odot}$ binary having optimal SNR $\sim 103$ corresponding to $D_{L}=100\,{\rm Mpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 250\,R_{s}$.
  2. 2.
    With ET, we can successfully extract the mass of the same SMBH and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 6610$ corresponding to $D_{L}=100\,{\rm Mpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 500\,R_{s}$, and within a precision of $\sim 20\%$ up to $R\sim 200\,R_{s}$.
  3. 3.
    We can successfully infer the masses of SMBHs as massive as $10^{8}\,M_{\odot}$, and localise a $100-100\,M_{\odot}$ binary having optimal SNRs $\sim 7624$ and $\sim 10$, corresponding to $D_{L}=1{\rm Gpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 5\times 10^{4}\,R_{s}$ with DECIGO and LISA.
Note that the chosen value of $x$ is a random draw from a uniform distribution between $\frac{1}{1-e},\frac{1}{1+e}$. We also find that, for certain portions of the parameter space, we can achieve sub-percent precision in DECIGO (see lower left regions of the parameter space in Figure 3).
For binaries in NSCs with a Bahcall-Wolf (BW) potential profile, having a massive black hole (MBH) at their center, the kinematic parameters projected along the LOS will depend on $M_{\rm MBH}$, $r$, steepness of the potential profile ($\alpha$), radial ($v_{r}$) and angular ($\Omega$) speeds of the CBC’s CoM at coalescence, and $\cos\vartheta$. Hence, together with LOSA, LOSJ, LOSS, and LOSC, we will need two more derivatives of LOSV to break all degeneracies among the inferred parameters.
Figure 10 shows the relative errors in the measurements of $M_{\rm MBH}$ (top panel), $r$ (middle panel), and $\alpha$ (bottom panel) over a grid of MBH masses and the distances of the merger. We find that:
  1. 1.
    With DECIGO, we can extract the mass of a $10^{6}\,M_{\odot}$ MBH within a precision of $\sim 10\%$ together with $\alpha$, and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 2794$ corresponding to $D_{L}=1\,{\rm Gpc}$, with a sub-percent precision out to $\sim 3\times 10^{5}\,R_{s}$.
  2. 2.
    With LISA, we can extract the mass of a $10^{6}\,M_{\odot}$ MBH within a precision of $\sim 25\%$ together with $\alpha$ and localise a $100-100\,M_{\odot}$ binary having optimal SNR $\sim 10$ corresponding to $D_{L}=1\,{\rm Gpc}$, with a sub-percent precision out to $\sim 4\times 10^{4}\,R_{s}$.
It is clear from Figure 10 that we can extract $M_{\rm MBH}$ and $r$ with sub-percent precision for certain parts of this parameter space, for both DECIGO and LISA. However, we cannot do so for $\alpha$ with LISA, for the fiducial CBC we consider. This can be understood as follows. Larger magnitudes of kinematic parameters increase our ability to constrain environment parameters. Getting closer to the MBH does indeed increase these magnitudes; however, the MBH potential starts to dominate over the potential with exponent $\alpha$. As a result, the precision of $\alpha$ gets suppressed. There is therefore a sweet spot in the $M_{\rm MBH}-r$ parameter space where the precision on the inference of $\alpha$ is optimized. This spot changes, depending on intrinsic parameters of the CBC, including its orbital parameters [^8] Figure 10 does not clearly show this sweet spot because it lies within the highly-relativistic domain where our analysis is not carried out..
Figure 4: Effective Keplerian outer-orbit in an NSC with Bahcall-Wolf profile: The precision of the extraction of parameters over a grid of MBH masses $M_{\rm MBH}$ and merger distance $r$ for a binary system of component masses (from left to right): $30-30M_{\odot}$ and $100-100M_{\odot}$ at $1{\rm Gpc}$ in DECIGO and LISA, respectively, for $\alpha=1.75$ and $\cos\vartheta=0.9$ with radial and tangential velocities: $v_{r}=250{\rm km/s}$ and $r\Omega=100{\rm km/s}$ [^10] These numbers are consistent with observed proper motions of stars in the Galactic NSC (see Figure 6 of Schödel et al. (2009))., respectively. Top Panel: Relative error in the mass of the MBH. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm MBH}}{M_{\rm MBH}}$. Middle Panel: Relative error in the distance of the binary from the NSC’s center. The ‘dashed’ lines are the lines of constant $\frac{\Delta r}{r}$. Bottom Panel: Relative error in the steepness parameter $\alpha$. The ‘dotted’ lines are the lines of constant $\frac{\Delta\alpha}{\alpha}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
For binaries in GCs with a Plummer potential Plummer (1911); Aarseth et al. (1974) profile, kinematic parameters along our LOS will depend on the total mass of the cluster ($M_{0}$), $r$, Plummer radius ($a_{\mathrm{p}}$), $v_{r}$, $\Omega$, and $\cos\vartheta$. Therefore, similar to the BW case, we will need six derivatives of LOSV to uniquely determine parameters. However, the Plummer profile being shallower than the BW profile, recovering LOSC and higher time derivatives of LOSV is not possible for the range of parameters we consider. Hence, we only evaluate the LOSA, LOSJ, and LOSS (see Tiwari et al. (2024a) for the expressions) when constructing the GW phase modulations.
Figure 5 shows the relative errors in the measurements of $M_{0}$ (top panel), $r$ (middle panel), and $a_{\mathrm{p}}$ (bottom panel) over a grid of $M_{0}$ and $r$ for $a_{\rm p}=0.1\,{\rm pc}$. We find that with DECIGO, we can successfully extract $M_{0}$, $a_{\rm p}$, and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 27940$ corresponding to $D_{L}=100\,\,{\rm Mpc}$, for total cluster masses greater than $\sim 3\times 10^{5}\,M_{\odot}$. Even though we can successfully constrain $M_{0}$, $r$, and $a_{\mathrm{p}}$ out to $r\sim 0.4\mathrm{pc}$, it still requires the cluster to be as heavy as $10^{7}\,M_{\odot}$. Note that with LISA, LOSJ is measurable for a very limited portion of the parameter space we consider. Therefore, we do not show the corresponding results.
Figure 5: Effective Keplerian outer-orbit in a GC with Plummer profile: The precision of the extraction of parameters over a grid of cluster masses $M_{0}$ and merger distance $r$ for a binary system of component masses $30-30\,M_{\odot}$ at $100\mathrm{Mpc}$ in DECIGO with $\cos\vartheta=0.25$. Left Panel: Relative error in the total mass of the cluster. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{0}}{M_{0}}$. Middle Panel: Relative error in the distance of the binary from the GC’s center. The ‘dashed’ lines are the lines of constant $\frac{\Delta r}{r}$. Right Panel: Relative error in the Plummer radius where $a_{\rm p}=0.1{\rm pc}$. The ‘dotted’ lines are the lines of constant $\frac{\Delta a_{\mathrm{p}}}{a_{\mathrm{p}}}$. The pink regions and the arrows have the same meaning as in Figure 2.
Discussion.— The astrophysical processes that govern compact binary coalescences (CBCs) are at best partially understood, if at all Spera et al. (2022). Several models of the environments in which CBCs form, evolve, and merge have been proposed Mapelli (2021), although, there is no conclusive evidence to suggest that one formation channel is preferred over others. Indeed, some works suggest that multiple formation channels have contributed to the production of the detected stellar-mass CBCs to date Zevin et al. (2021); Pierra et al. (2024). The intrinsic parameters of a CBC, viz., its component masses, spins, and/or orbital eccentricity, can sometimes be suggestive of a subset of possible merger environments and formation channels Farr et al. (2017); Bavera et al. (2020); Bogdanovic et al. (2007); Vajpeyi et al. (2022); McKernan and Ford (2024); Adamcewicz et al. (2024); Samsing et al. (2022); Tagawa et al. (2021); Kozai (1962); Lidov (1962); Naoz (2016); Antonini et al. (2017); Yu et al. (2020); Randall and Xianyu (2018); Gerosa and Fishbach (2021); Liu and Lai (2021). They cannot, however, always conclusively rule out all other possible formation channels, especially if there is no EM counterpart.
In this work we show for the first time that, on a single-event basis, rich and precise information of the environment with generic gravitational potentials can be extracted. Additionally, we show that it also possible to acquire information about the location of the CBC with respect to the center of the potential. This can be achieved provided the CBC’s LOSA and some of its higher time derivatives can be measured. We consider the SMBH, NSC (BW), and GC (Plummer) potentials. We find that we can measure enclosed mass and potential profile exponents, and the eccentricity and the radius of the CBC’s orbit in the potential, with relative errors of less than unity for a large portion of the parameter space we consider, even going down to sub-percent precision for certain portions.
We expect that our work, demonstrating the possibility of constraining environment parameters, on a single-event basis, could have significant astrophysical implications, potentially answering the longstanding question of the provenance of CBCs for at least a fraction of them with measurable LOSV time derivatives [^11] See Tiwari et al. (2024b) for estimates of this fraction (focusing exclusively on LOSAs) for various GC configurations.. Moreover, it could unearth important correlations between the mass of the CBC, the mass of the environment, the steepness of the potential profile, and the location of the CBC with respect to the center of the potential. Probing these would directly test astrophysical mechanisms, such as dynamical friction and mass segregation, among others, in dense stellar environments.
Strictly, our work probes the environment near the end of the CBC’s lifetime. Given that the delay time between the CBC’s formation and merger could span millions to billions of years (see, e.g., Fishbach and Kalogera (2021), and references therein), the location of the environment probed by the CBC could be quite different from the location of its formation, due to kicks induced by, e.g., asymmetric supernova mechanisms, or interactions with other bodies [^12] The probability of such interactions are relatively higher in dense star clusters such as NSCs.. Thus, our work complemented by EM observations of stellar binaries could enable the construction of the complete story of a binary’s evolution, starting out as well-separated (bound) main-sequence stars and ending with the coalescence of compact objects.
Acknowledgements.
We thank Debatri Chattopadhyay for their careful reading and feedback. We thank K. G. Arun, Debarati Chatterjee and Zoheyr Doctor for useful discussions. SJK gratefully acknowledges support from SERB Grants SRG/2023/000419 and MTR/2023/000086. SC acknowledges support from the Department of Atomic Energy, Government of India, under project no. 12-R&D-TFR-5.02-0200 and RTI 4002. All computations were performed on the Sarathi Computing Cluster at IUCAA.
Software: NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2020), astropy (Robitaille et al., 2013; Price-Whelan et al., 2018), Matplotlib (Hunter, 2007), jupyter (Kluyver et al., 2016), Numdifftools (Numdifftools, 2014).